Consider the two-port network shown with n=3, R=9k and C=6?F.

(50 points) Consider the two-port network shown with n - 3, R - 9kS2 and C -6pF. a. Determine the hybrid [h] and impedance [z] parameters (as functions of s; hint: each hybrid parameter is a first order transfer function) 1000 ? 10.333333 0.3333 6.667E-17 166666.676 | ? | 500000/s 500000/s |?| 9000/s
Answer #1

(a) Consider the circuit. 1:3 9 k22 76 LF 3 & c a The voltage across the capacitor is equal to the input voltage, V. V, =(9 k

Consider the equations for h-parameters. V = h,1 +h2V2 12 = h, 1+hV, VA 12 11 11,50 V, = 0 represents a short circuit at the

12 = 3x10-6 sv. 12 =(3x10s)[V> –(9 k)12] (1+0.027s)), = 3x10-65V, h - 123x1065 V2 Ip-0 (1+0.0275) h, _ 1.11x104 S+ 37.04 (b)